what number is 120% of 80?
how do i do this?

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

what number is 120% of 80?
how do i do this?
-----------------
That % sign means "over 100"
So, 120% means 120/100 which is 1.2
Your Problem:
(120%)*80 = 1.2*80 = 96
===============
Cheers,
Stan H.

Answer by checkley71(8403)

120%=1.2
1.2*80=96 ANSWER.